# Aditya Tanna, Pratinav Seth, Mohamed Bouadi Utsav Avaiya, Vinay Kumar Sankarapu Lexsi Labs, India & France

### v1

@@ 1  Introduction
Traditional approaches like gradient-boosted decision trees (e.g., XGBoost [1], LightGBM [2], CatBoost [3]) remain dominant because they reliably handle heterogeneous features, variable sample sizes, and small datasets—challenges that standard neural networks often struggle with. Deep models frequently underperform on tabular benchmarks due to the lack of spatial or temporal structure, making feature extraction less amenable to convolutional or sequential architectures. Moreover, tabular datasets often exhibit high heterogeneity—mixing numerical, categorical, and ordinal variables—and suffer from missing values and skewed distributions, which violate many of the inductive biases that make deep learning successful in other modalities. As a result, practitioners continue to favor tree ensembles and gradient boosting for their robustness, interpretability, and low data requirements.
Recently, the advent of tabular foundation models (TFMs) has begun to reshape this landscape. These models extend the “foundation model” paradigm—pretraining on large, diverse corpora followed by downstream adaptation—to structured data domains. Models such as TabPFN [4, 5], TabICL [6], OrionMSP [7], and OrionBiX [8] leverage large-scale pre-training on synthetic or massive tabular datasets to provide generic, out-of-the-box learning capabilities.
@@
  • A systematic benchmarking module across standard tabular suites (e.g., TALENT [9], OpenML-CC18 [10]) that evaluates accuracy, calibration, fairness, and resource efficiency in a unified ranking framework, facilitating reproducible and comparative studies of TFM performance.
@@ 2.1  Traditional Machine Learning and Deep Learning for Tabular Data
In recent years, deep neural networks—such as feed-forward multilayer perceptrons, attention-based architectures, and hybrid tree–neural models—have been applied to tabular data [11, 12]. Despite significant progress, these models have often struggled to consistently outperform GBDTs in real-world benchmarks. This gap arises from several challenges: the heterogeneity of tabular data, limited dataset sizes, complex feature–target relationships, and the absence of large-scale pre-training strategies analogous to those that revolutionized vision and language modeling.
@@ 2.2  Tabular Foundation Models
TabPFN (Tabular Prior-data Fitted Network) represents one of the earliest and most influential TFMs. It employs a transformer architecture trained on vast collections of synthetic tabular datasets, effectively learning to approximate Bayesian inference across varying generative priors. This pre-training strategy enables TabPFN to achieve strong few-shot and zero-shot performance on small to medium datasets (typically up to $\sim$10,000 samples), offering competitive accuracy without explicit fine-tuning [4, 5].
@@
Building on this foundation, TabICL extends the in-context learning paradigm to larger-scale tabular datasets by introducing a novel two-stage column-then-row attention mechanism. Through large-scale synthetic pre-training (up to $\sim$60K samples) and robust scaling to half a million samples at inference time, TabICL demonstrates that transformer-based ICL can challenge traditional tree ensembles even on large tabular domains [6].
@@
Another direction, TabDPT, introduces denoising-based or self-supervised pre-training strategies that learn robust feature representations from both synthetic and real-world tables [13]. In parallel, other approaches explore the role of mixed synthetic priors, showing that a combination of causal, statistical, and randomized priors during pre-training can substantially improve downstream generalization and calibration performance across classification and regression tasks. Collectively, these developments illustrate that, with sufficient architectural bias and large-scale pre-training, transformer-based TFMs can achieve broad generalization on structured data—challenging long-held assumptions about the superiority of tree-based ensembles in tabular learning. Recent comprehensive surveys have catalogued the landscape of tabular foundation models [14, 15], highlighting their potential to transform tabular learning. Recent work has demonstrated that careful fine-tuning strategies can yield significant improvements for TFMs [16]. Meanwhile, researchers have argued that tabular foundation models should be a research priority [17], given their potential impact across domains.
@@
OrionMSP [7] introduces a transformer architecture for tabular in-context learning that combines multi-scale sparse attention—capturing both local and global feature dependencies efficiently—with a Perceiver-style latent memory that enables safe bidirectional communication between training and test samples. It uses column-wise embeddings via Set Transformers to model feature distributions and a split-masked causal attention mechanism for proper ICL reasoning. The design achieves near-linear attention scaling while maintaining high accuracy across diverse datasets, rivaling or outperforming TabICL and TabPFN—especially on heterogeneous, high-dimensional, and imbalanced datasets—thus demonstrating robust, efficient, and generalizable tabular learning without gradient updates.
@@
Similarly, OrionBiX [8] employs a biaxial attention mechanism for tabular in-context learning, enabling bidirectional context modeling across both rows and columns of tabular data. This dual-axis attention pattern allows the model to capture complex feature interactions and sample relationships simultaneously, providing a more comprehensive representation of tabular structure. The biaxial design facilitates effective information flow between training and test samples, enabling strong in-context learning performance across diverse tabular datasets while maintaining computational efficiency.
@@ 2.3  Representation Learning and Alternative Approaches
While transformer-based TFMs dominate recent work, alternative representation learning paradigms have been explored. Graph neural networks have been adapted for tabular data, modeling feature relationships through graph structures [18]. Extensions of foundation models to time series forecasting demonstrate their versatility; for instance, TabPFN-v2 has shown competitive performance on temporal tasks [19], suggesting that tabular foundation models may generalize beyond traditional classification settings.
@@ 2.4  Tabular ML Toolkits and Frameworks
Several frameworks have emerged to facilitate deep learning with tabular data, each addressing different aspects of the modeling pipeline. AutoGluon [20] is an AutoML framework that automates model selection and hyperparameter tuning for tabular data, employing neural architecture search and ensemble methods to achieve strong predictive performance with minimal manual configuration. PyTorch Tabular [21] provides a modular framework for deep learning with tabular data, offering reusable components for neural network architectures tailored to structured data. However, these frameworks primarily focus on training models from scratch or performing architecture search, rather than working with pretrained tabular foundation models.
@@ 2.5  Adaptation Techniques: Meta-Learning, Fine-Tuning, and PEFT
Meta-learning (or episodic fine-tuning) uses support–query splits from training data to retain in-context generalization while improving task-specific performance [22]. This approach has proven effective for tabular foundation models such as TabICL, where episodic training enhances accuracy and supports few-shot adaptation to new distributions [6, 13].
Supervised fine-tuning (SFT) updates all model parameters on labeled data when sufficient resources are available [23]. In tabular domains, full fine-tuning often yields notable gains over inference-only use, especially on large datasets where models learn domain-specific patterns [6, 13]. However, it remains computationally and storage intensive, as every parameter must be updated per task.
Parameter-efficient fine-tuning (PEFT) methods, such as Low-Rank Adaptation (LoRA) [24], limit updates to low-dimensional subspaces, greatly reducing compute and memory costs while preserving performance. Surveys categorize PEFT into additive (adapters), selective (partial updates), and reparameterization-based (low-rank) approaches [25]. Empirical studies show that PEFT matches or surpasses full fine-tuning with far lower resource demands [26]. For tabular foundation models, PEFT enables efficient domain adaptation under tight resource budgets, supporting rapid experimentation and scalable deployment. Together, these paradigms bridge the gap between zero-shot inference and fully supervised modeling, promoting adaptable TFMs for diverse applications.
@@ 2.6  Calibration, Fairness, and Responsible Deployment
Calibration reliability—the alignment between predicted probabilities and actual outcomes—is vital for high-stakes decision-making. Guo et al. [27] showed that modern neural networks often exhibit systematic miscalibration, motivating metrics such as Expected Calibration Error (ECE) and Maximum Calibration Error (MCE) to measure these gaps. For tabular foundation models, early results suggest that pretrained transformers yield better-calibrated confidence estimates than classical ensembles, especially in low-data settings where in-context learning supports more nuanced uncertainty estimation.
Fairness requires that models avoid systematic bias across demographic groups [28]. Fairness frameworks propose criteria such as equalized odds [29]—requiring equal true and false positive rates—and equal opportunity, focusing on true positive rate parity. [30] revealed a fundamental tension between perfect calibration and strict fairness, showing that these objectives can conflict. In-context learning setups may further amplify biases present in demonstration data. Mitigation strategies include group-balanced sampling, fair demonstration selection, and uncertainty-based data curation.
@@ 4.1  Model-Aware Preprocessing
The DataProcessor dynamically loads model-specific routines to format data correctly with minimal user setup. For TabPFN, features are mapped to numerical or integer-encoded tensors to align with its synthetic priors. ContextTab [31] adds semantic vectors for column names and categorical values. TabICL and Mitra apply normalization and model-tailored transformations. For OrionMSP [7] and OrionBiX [8], preprocessing is distribution- and hierarchy-aware: columns are embedded with Set Transformer-based induced set attention on train rows only, numerical features are standardized and categorical values embedded, multi-scale row inputs and block-sparse attention masks are constructed, Perceiver-style memory is written by train rows and read by all rows, and label injection uses split-masked attention to prevent test-to-train leakage. Compared to TabICL [6], Orion adds distribution-aware column embeddings and multi-scale sparsity for scalable, ICL-safe processing. Together, these routines ensure each TFM receives inputs in its expected schema.
@@ 6.1  Experimental Design
Our experiments encompass three well-established benchmark suites—TALENT [9] (181 automatically discovered datasets), OpenML-CC18 [10] (72 curated datasets), and TabZilla [32] (36 heterogeneous tasks)—facilitating a systematic comparison across a broad spectrum of tabular learning settings. We further conduct domain-specific evaluations in high-impact areas such as medical and financial domains to assess the practical applicability of the methods. All experiments adhere to the standardized dataset splits defined by each benchmark, ensuring both reproducibility and fairness in comparison.
@@
We evaluate seven recent tabular foundation models—TabPFN, TabICL, OrionMSP, OrionBiX, Mitra, ContextTab, and TabDPT—under multiple adaptation paradigms, including zero-shot inference, meta-learning fine-tuning, supervised fine-tuning, parameter-efficient supervised fine-tuning, and parameter-efficient meta-learning. In addition, we include established traditional baselines using autogloun [20] such as XGBoost, LightGBM, CatBoost, and Random Forest as strong reference models for comparison.

### v2

@@ 1  Introduction
Traditional approaches like gradient-boosted decision trees (e.g., XGBoost [chen2016xgboost], LightGBM [ke2017lightgbm], CatBoost [prokhorenkova2018catboost]) remain dominant because they reliably handle heterogeneous features, variable sample sizes, and small datasets—challenges that standard neural networks often struggle with. Deep models frequently underperform on tabular benchmarks due to the lack of spatial or temporal structure, making feature extraction less amenable to convolutional or sequential architectures. Moreover, tabular datasets often exhibit high heterogeneity—mixing numerical, categorical, and ordinal variables—and suffer from missing values and skewed distributions, which violate many of the inductive biases that make deep learning successful in other modalities. As a result, practitioners continue to favor tree ensembles and gradient boosting for their robustness, interpretability, and low data requirements.
Recently, the advent of tabular foundation models (TFMs) has begun to reshape this landscape. These models extend the “foundation model” paradigm—pretraining on large, diverse corpora followed by downstream adaptation—to structured data domains. Models such as TabPFN [tabpfn, TabPFN2], TabICL [tabicl], OrionMSP [orionmsp2025], and OrionBiX [orionbix2025] leverage large-scale pre-training on synthetic or massive tabular datasets to provide generic, out-of-the-box learning capabilities.
@@
  • A systematic benchmarking module across standard tabular suites (e.g., TALENT [talent], OpenML-CC18 [openmlcc18]) that evaluates accuracy, calibration, fairness, and resource efficiency in a unified ranking framework, facilitating reproducible and comparative studies of TFM performance.
@@ 2.1  Traditional Machine Learning and Deep Learning for Tabular Data
In recent years, deep neural networks—such as feed-forward multilayer perceptrons, attention-based architectures, and hybrid tree–neural models—have been applied to tabular data [borisov2022tabular, somvanshi2024tabular]. Despite significant progress, these models have often struggled to consistently outperform GBDTs in real-world benchmarks. This gap arises from several challenges: the heterogeneity of tabular data, limited dataset sizes, complex feature–target relationships, and the absence of large-scale pre-training strategies analogous to those that revolutionized vision and language modeling.
@@ 2.2  Tabular Foundation Models
TabPFN (Tabular Prior-data Fitted Network) represents one of the earliest and most influential TFMs. It employs a transformer architecture trained on vast collections of synthetic tabular datasets, effectively learning to approximate Bayesian inference across varying generative priors. This pre-training strategy enables TabPFN to achieve strong few-shot and zero-shot performance on small to medium datasets (typically up to $\sim$10,000 samples), offering competitive accuracy without explicit fine-tuning [tabpfn, TabPFN2].
@@
Building on this foundation, TabICL extends the in-context learning paradigm to larger-scale tabular datasets by introducing a novel two-stage column-then-row attention mechanism. Through large-scale synthetic pre-training (up to $\sim$60K samples) and robust scaling to half a million samples at inference time, TabICL demonstrates that transformer-based ICL can challenge traditional tree ensembles even on large tabular domains [tabicl].
@@
Another direction, TabDPT, introduces denoising-based or self-supervised pre-training strategies that learn robust feature representations from both synthetic and real-world tables [tabdpt]. In parallel, other approaches explore the role of mixed synthetic priors, showing that a combination of causal, statistical, and randomized priors during pre-training can substantially improve downstream generalization and calibration performance across classification and regression tasks. Collectively, these developments illustrate that, with sufficient architectural bias and large-scale pre-training, transformer-based TFMs can achieve broad generalization on structured data—challenging long-held assumptions about the superiority of tree-based ensembles in tabular learning. Recent comprehensive surveys have catalogued the landscape of tabular foundation models [jiang2025representation, badaro2023transformers], highlighting their potential to transform tabular learning. Recent work has demonstrated that careful fine-tuning strategies can yield significant improvements for TFMs [rubachev2025finetuning]. Meanwhile, researchers have argued that tabular foundation models should be a research priority [breugel2024position], given their potential impact across domains.
@@
OrionMSP [orionmsp2025] introduces a transformer architecture for tabular in-context learning that combines multi-scale sparse attention—capturing both local and global feature dependencies efficiently—with a Perceiver-style latent memory that enables safe bidirectional communication between training and test samples. It uses column-wise embeddings via Set Transformers to model feature distributions and a split-masked causal attention mechanism for proper ICL reasoning. The design achieves near-linear attention scaling while maintaining high accuracy across diverse datasets, rivaling or outperforming TabICL and TabPFN—especially on heterogeneous, high-dimensional, and imbalanced datasets—thus demonstrating robust, efficient, and generalizable tabular learning without gradient updates.
@@
Similarly, OrionBiX [orionbix2025] employs a biaxial attention mechanism for tabular in-context learning, enabling bidirectional context modeling across both rows and columns of tabular data. This dual-axis attention pattern allows the model to capture complex feature interactions and sample relationships simultaneously, providing a more comprehensive representation of tabular structure. The biaxial design facilitates effective information flow between training and test samples, enabling strong in-context learning performance across diverse tabular datasets while maintaining computational efficiency.
@@ 2.3  Representation Learning and Alternative Approaches
While transformer-based TFMs dominate recent work, alternative representation learning paradigms have been explored. Graph neural networks have been adapted for tabular data, modeling feature relationships through graph structures [li2024gnn]. Extensions of foundation models to time series forecasting demonstrate their versatility; for instance, TabPFN-v2 has shown competitive performance on temporal tasks [hoo2025tables], suggesting that tabular foundation models may generalize beyond traditional classification settings.
@@ 2.4  Tabular ML Toolkits and Frameworks
Several frameworks have emerged to facilitate deep learning with tabular data, each addressing different aspects of the modeling pipeline. AutoGluon [erickson2020autogluon] is an AutoML framework that automates model selection and hyperparameter tuning for tabular data, employing neural architecture search and ensemble methods to achieve strong predictive performance with minimal manual configuration. PyTorch Tabular [joseph2021pytorch] provides a modular framework for deep learning with tabular data, offering reusable components for neural network architectures tailored to structured data. However, these frameworks primarily focus on training models from scratch or performing architecture search, rather than working with pretrained tabular foundation models.
@@ 2.5  Adaptation Techniques: Meta-Learning, Fine-Tuning, and PEFT
Meta-learning (or episodic fine-tuning) uses support–query splits from training data to retain in-context generalization while improving task-specific performance [finn2017model]. This approach has proven effective for tabular foundation models such as TabICL, where episodic training enhances accuracy and supports few-shot adaptation to new distributions [tabicl, tabdpt].
Supervised fine-tuning (SFT) updates all model parameters on labeled data when sufficient resources are available [howard2018universal]. In tabular domains, full fine-tuning often yields notable gains over inference-only use, especially on large datasets where models learn domain-specific patterns [tabicl, tabdpt]. However, it remains computationally and storage intensive, as every parameter must be updated per task.
Parameter-efficient fine-tuning (PEFT) methods, such as Low-Rank Adaptation (LoRA) [hu2021lora], limit updates to low-dimensional subspaces, greatly reducing compute and memory costs while preserving performance. Surveys categorize PEFT into additive (adapters), selective (partial updates), and reparameterization-based (low-rank) approaches [liu2024peft]. Empirical studies show that PEFT matches or surpasses full fine-tuning with far lower resource demands [razuvayevskaya2024peft]. For tabular foundation models, PEFT enables efficient domain adaptation under tight resource budgets, supporting rapid experimentation and scalable deployment. Together, these paradigms bridge the gap between zero-shot inference and fully supervised modeling, promoting adaptable TFMs for diverse applications.
@@ 2.6  Calibration, Fairness, and Responsible Deployment
Calibration reliability—the alignment between predicted probabilities and actual outcomes—is vital for high-stakes decision-making. Guo et al. [guo2017calibration] showed that modern neural networks often exhibit systematic miscalibration, motivating metrics such as Expected Calibration Error (ECE) and Maximum Calibration Error (MCE) to measure these gaps. For tabular foundation models, early results suggest that pretrained transformers yield better-calibrated confidence estimates than classical ensembles, especially in low-data settings where in-context learning supports more nuanced uncertainty estimation.
Fairness requires that models avoid systematic bias across demographic groups [barocas2019fairness]. Fairness frameworks propose criteria such as equalized odds [hardt2016equality]—requiring equal true and false positive rates—and equal opportunity, focusing on true positive rate parity. [pleiss2017fairness] revealed a fundamental tension between perfect calibration and strict fairness, showing that these objectives can conflict. In-context learning setups may further amplify biases present in demonstration data. Mitigation strategies include group-balanced sampling, fair demonstration selection, and uncertainty-based data curation.
@@ 4.1  Model-Aware Preprocessing
The DataProcessor dynamically loads model-specific routines to format data correctly with minimal user setup. For TabPFN, features are mapped to numerical or integer-encoded tensors to align with its synthetic priors. ContextTab [contexttab] adds semantic vectors for column names and categorical values. TabICL and Mitra apply normalization and model-tailored transformations. For OrionMSP [orionmsp2025] and OrionBiX [orionbix2025], preprocessing is distribution- and hierarchy-aware: columns are embedded with Set Transformer-based induced set attention on train rows only, numerical features are standardized and categorical values embedded, multi-scale row inputs and block-sparse attention masks are constructed, Perceiver-style memory is written by train rows and read by all rows, and label injection uses split-masked attention to prevent test-to-train leakage. Compared to TabICL [tabicl], Orion adds distribution-aware column embeddings and multi-scale sparsity for scalable, ICL-safe processing. Together, these routines ensure each TFM receives inputs in its expected schema.
@@ 6.1  Experimental Design
Our experiments encompass three well-established benchmark suites—TALENT [talent] (181 automatically discovered datasets), OpenML-CC18 [openmlcc18] (72 curated datasets), and TabZilla [tabzilla] (36 heterogeneous tasks)—facilitating a systematic comparison across a broad spectrum of tabular learning settings. We further conduct domain-specific evaluations in high-impact areas such as medical and financial domains to assess the practical applicability of the methods. All experiments adhere to the standardized dataset splits defined by each benchmark, ensuring both reproducibility and fairness in comparison.
@@
We evaluate seven recent tabular foundation models—TabPFN, TabICL, OrionMSP, OrionBiX, Mitra, ContextTab, and TabDPT—under multiple adaptation paradigms, including zero-shot inference, meta-learning fine-tuning, supervised fine-tuning, parameter-efficient supervised fine-tuning, and parameter-efficient meta-learning. In addition, we include established traditional baselines using autogloun [erickson2020autogluon] such as XGBoost, LightGBM, CatBoost, and Random Forest as strong reference models for comparison.